OpenStudy (anonymous):

Solve the following by subtracting the exponents. 2. s3t4 ______ s2t2 3. q3r2s _____ q2rs

OpenStudy (lalaly):

\[\huge{s^{3-2}t^{4-2}} =st^2\]

OpenStudy (lalaly):

why dont you try the second one on ur own:)

OpenStudy (anonymous):

qr2s

OpenStudy (anonymous):

that the answer i had

OpenStudy (lalaly):

no \[\huge{q^{3-2}r^{2-1}s^{1-1}}= qr\]

OpenStudy (anonymous):

wow ok im learning thanks

OpenStudy (lalaly):

:)